# Via Pietro Giuria 1, I-10125 Turin, Italy

### v1

@@ Abstract
The study of entanglement in gauge theories is expected to provide insights into many fundamental phenomena, including confinement. However, calculations of quantities related to entanglement in gauge theories are limited by ambiguities that stem from the non-factorizability of the Hilbert space. In this work we study lattice gauge theories that admit a dual description in terms of spin models, for which the replica trick and Rényi entropies are well defined. In the first part of this work, we explicitly perform the duality transformation in a replica geometry, deriving the structure of a replica space for a gauge theory. Then, in the second part, we calculate, by means of Monte Carlo simulations, the entropic c-function of the $\mathbb{Z}_{2}$ gauge theory in three spacetime dimension, exploiting its dual description in terms of the three-dimensional Ising model.
@@ 1 Introduction
In quantum field theories the entanglement entropy is affected by ultraviolet divergences: for highly excited states the divergence is proportional to the spatial volume of the system, whereas ground states exhibit the so-called “area law” Eisert:2008ur , namely the ultraviolet divergence is proportional to the area of the entangling surface $\partial A$ separating $A$ and $B$. Therefore, a proper regularization is required, in order to identify the subleading contributions, which encode important physical information about the “resolution scale” at which the degrees of freedom of the system are probed Casini:2010kt ; Liu:2012eea ; Casini:2012ei ; Myers:2012ed ; Klebanov:2012va ; Liu:2013una ; Jokela:2024cxb , the geometry of the system Klebanov:2012yf ; Casini:2014yca ; Mezei:2014zla ; Faulkner:2015csl ; Bianchi:2016xvf and the nature of the state itself Casini:2008cr . For a system defined in $D=d+1$ spacetime dimensions, it is then convenient to consider a geometry in which the entangling surface does not depend on the size of the subsystem $A$; a common choice is the slab geometry, in which $A$ is maximally extended in all space-like directions except for the one separating $A$ and $B$, where it has a finite size $l$. One can then define the ultraviolet-finite entropic (Rényi) c-function as Nishioka:2006gr ; Casini:2006es ; Casini:2004bw
@@
Calculations of such highly non-local quantities are typically challenging, both analytically and numerically. A tool which is commonly used is the replica trick Calabrese:2004eu ; Calabrese:2009qy , which allows to express the trace of a power of the reduced density matrix appearing in eq. (1) in terms of ratios of partition functions,
@@
A non-trivial aspect in the study of entanglement entropy in gauge theories, however, is the ambiguity of the definition itself in the presence of a local symmetry. Gausß’s law and consequently the non-local constraint defining physical states makes the physical Hilbert space of a gauge theory non-factorizable. This problem has been extensively discussed in the literature, and different definitions of entanglement entropy have been proposed Buividovich:2008gq ; Donnelly:2011hn ; Agon:2013iva ; Casini:2013rba ; Radicevic:2014kqa ; Ohmori:2014eia ; Donnelly:2014gva ; Ghosh:2015iwa ; Aoki:2015bsa ; Chen:2015kfa ; Radicevic:2015sza ; Soni:2015yga ; VanAcoleyen:2015ccp ; Radicevic:2016tlt ; Soni:2016ogt ; Aoki:2016lma ; Lin:2018bud , among which two approaches are most commonly used: one can either extend the Hilbert space of states, including gauge non-invariant ones to ensure factorizability; or work at the level of the algebra of observables defined over a specific region, specifying which generators have to be included in region $A$ and which ones in its complement. It turns out that for some choices of the algebras the two approaches are equivalent Soni:2015yga , and lead to the following general result: the trace which defines the entanglement entropy splits into the sum of different contributions, labeled by the eigenvalues of the operators at the boundary between the system $A$ and its complement, which in general belong to the center of the algebra of observables of $A$. As a result, in Abelian gauge theories defined on the lattice the entanglement entropy admits the following decomposition
@@
In this work we address one facet of this problem, focusing on the construction of a replica space for gauge theories through a duality transformation. For gauge theories defined in three spacetime dimensions, the latter can be carried out through a Kramers–Wannier transform Kramers:1941kn ; Kramers:1941zz , which maps the gauge theory to a spin system. It is well known that such dualization is not straightforward when the gauge theory is defined on a space with a boundary and there exist operators (including, in particular, those defining Gausß’s law) defined on the boundary. However, it was shown in different works Casini:2014aia ; Radicevic:2016tlt ; Moitra:2018lxn ; Lin:2018bud that, even though the entanglement entropy is not preserved by the duality transformation, its universal part, encoded in the entropic c-function, is exactly mapped in the continuum limit.
In the first part of this work, we explicitly perform the duality transformation starting from a spin model defined on a lattice in three spacetime dimensions with an arbitrary positive integer number of replicas, to derive the corresponding geometry for the dual gauge theory. As we discuss in what follows, the physical results we obtain from our construction are purely geometric, namely they do not depend on the degrees of freedom of the theory (nor on the details of how they are assigned to $A$ or $B$); rather, the duality transformation gives us a well-defined prescription on how to implement a conical singularity on the lattice. The logical scheme of how the direct and the dual formulation of a theory are connected with each other, in the Hamiltonian and in the Lagrangian formalisms, is summarized in fig. 1 (which also includes the references to previous literature, in which the various aspects discussed herein were studied).
@@
The structure of the article is the following. In section 2 we review the generalities about the duality transformation, and discuss its formulation in the presence of replicas; we mostly focus on the case of the Ising model, considering both the case of two and three spacetime dimensions, but we also discuss the extension of our results to other gauge theories, including $\mathrm{U}(1)$ and $\mathbb{Z}_{N}$ gauge theories (and briefly comment on the generalization to higher dimensions and to non-Abelian gauge theories). Section 3 illustrates the connection between the Hamiltonian and the Lagrangian formulations in the presence of replicas, with a detailed discussion of the $(1+1)$-dimensional Ising model case study, while section 4 is devoted to a Monte Carlo study of the entropic c-function of the three-dimensional $\mathbb{Z}_{2}$ gauge theory. Finally, in section 5 we discuss our findings and comment on its possible generalizations. The appendix contains a discussion of the rôle of boundary conditions in the duality transformation (section A), a study of the dual replica space for other two-dimensional spin models (section B), a derivation of the spin theory dual to the three-dimensional $\mathrm{U}(1)$ gauge theory (section C), and some technical details about our numerical simulations (section D).
@@ 2.2 Duality in the presence of replicas
Figure 2: Two-dimensional replica space. The vertical direction is the Euclidean-time direction, while the horizontal one is the spatial one. The red wavy line represents the cut; green links connect spins in different replicas, therefore the periodicity of the lattice in the subsystem $A$ is $n$ times the periodicity in subsystem $B$.
@@
The geometry we are interested in is the one of $n$ replicas connected through a cut, and let us first consider the two-dimensional case, as sketched in fig. 2. In a replica geometry, the expression (8) still holds, since its derivation is only based on the nearest-neighbor nature of the interaction and is independent from the geometry of the lattice. The difference in the duality transformation is in the representation of the $k_{i,\mu}$ variables that satisfy the constraint; clearly, the introduction of a cut does not affect the duality transformation for links not crossed by the cut, since the transformation only involves nearest-neighbors spins in the dual lattice and the links in the direct lattice separating them. Moreover, the duality transformation remains unchanged also for links crossed by the cut that are far from the cut boundary. [^1] This is related to the fact that the exact location of the cut is irrelevant. The only difference arises at the edges of the cut; in particular, in two dimensions the dual lattice has sites exactly at (the lattice analogue of) the branch points of the replica space. Consider, for example, the cartoon in fig. 3, where two of the $n$ replicas are depicted: the equation expressing the constraint associated with the site $i$ of the direct lattice, in the replica labeled $1$, is
@@
Figure 4: Duality transformation in a replica space. The solid line is the entangling surface along the links of the dual lattice, the red region is the cut, the red dot the intersection between the link between the sites $i$ and $i+\hat{0}$ and the cut.
It is straightforward to carry out the dualization in the presence of replicas for the three-dimensional Ising model, too. In this case (see the sketch in fig. 4) the “cut” is actually a surface, and the conical singularity is a line where the cut ends. As for the two-dimensional model, the way the duality maps degrees of freedom in the direct and in the dual formulation is not affected by the presence of conical singularities, for spins far enough from them. Consider, for example, the site $i$ in the direct lattice shown in fig. 4; in the first replica, the Kronecker delta constraint is of the form
@@
The presence of gauge fields shared among different replicas has some constraints on the gauge transformations that leave the physical properties of the system invariant. In particular, there exist plaquette terms that can be schematically written as $U_{\mbox{\tiny{edge}}}U^{(r)}_{\mbox{\tiny{staple}}}$, where $U_{\mbox{\tiny{edge}}}$ is the gauge field on the conical singularity and $U^{(r)}_{\mbox{\tiny{staple}}}$ is a staple in replica $r$ forming a plaquette together with $U_{\mbox{\tiny{edge}}}$. The presence of such terms requires that gauge transformations performed on the sites along the entangling surface have to act on all replicas. More precisely, a gauge transform on a site on the entangling surface acts upon $(4n+2)$ gauge fields: the $4n$ links orthogonal to the conical singularity attached to that point, and the two links along the entangling surface shared by all replicas. This restriction on the possible gauge transformations in a replica space was also noted in the context of $\mathrm{SU}(N)$ gauge theories Soni:2016ogt ; Jokela:2022fvh .
@@
The difference between the Rényi entropies in the gauge and in the spin theories is the presence of a ultraviolet-divergent term proportional to the length of the entangling surface, therefore contributing to the area law. This term makes the Rényi entropy of the gauge theory larger than the one of the Ising model, as was already pointed out in the literature Tagliacozzo:2010vk . The origin of this difference stems from the constraint on the number of possible gauge transforms that can be defined in a system with replicas, and results into a multiplicative constant affecting the partition function for the dual gauge theory, which obviously does not carry any physical information. Indeed, one can also note that the term $|\partial A|-1$ depends on the gauge choice at the entangling surface, as was also discussed in refs. Casini:2013rba ; Aoki:2015bsa ; Aoki:2016lma . Like in the $D=2$ case, physical information can be obtained from a suitably defined entropic c-function, in which the area law term cancels out; in $D=3$ if the length of the entangling surface does not depend on the size of the region $A$, as in the slab geometry, the entropic c-function can be simply defined according to eq. (2), and one gets
@@
It is also worth noting that, in the opposite limit ($\beta\rightarrow\infty$, or $\beta^{\star}\rightarrow 0$) the result one would naïvely obtain is not the same as the Hamiltonian formalism:
@@
So far, we assumed the entangling surface to lie on the lattice dual to the Ising model, that is, to be placed along the links of the dual gauge theory. However, it is also interesting to analyze a geometry in which the entangling surface lies on the links of the spin model; in this case there are spin degrees of freedom on the conical singularity. From the discussion of the $D=2$ case, we know that such degrees of freedom are shared among all replicas, and have a larger number of nearest neighbors, which here is $(4n+2)$, of which $4n$ in the directions orthogonal to the entangling surface and $2$ along the conical singularity.
@@
where $U_{\Box\times n}$ denotes the closed loop of length $4n$ that winds around the conical singularity. Note that the presence of two such loops of length $4n$ is necessary to satisfy the Kronecker-delta constraint around a point of the direct lattice with an enhanced number of nearest neighbors: for the $4n$ links touching a site on the conical singularity one can make the usual duality transformation according to eq. (12), but then there are only two link variables to be adjusted, such that in the sum $\sum_{i}k_{i,\mu}$ all gauge variables appear twice. As a consequence, the two $k$ variables along the conical singularity must be built from a larger number of gauge variables. Note that this geometry was already discussed in ref. Chen:2015kfa , whose authors carried out a strong-coupling calculation of the Rényi entropy in this “central-plaquette” geometry, where the central plaquettes are the ones encircling the conical singularity.
@@
Figure 5: Central-plaquette geometry. The blue loop crosses the cut (red surface) and therefore has length $4n$, where $n$ is the number of replicas.
@@ 2.3 $\mathrm{U}(1)$ and $\mathbb{Z}_{N}$ models
In the presence of replicas, the duality transformation can be generalized along the lines that we discussed for the three-dimensional Ising model; the geometry we consider is the same as in fig. 4, but with $\phi$ variables instead of $\sigma$ ones. The constraint associated to the site $i$ of the left replica is again expressed by eq. (19). Note that, with the cut placed as in fig. 4, all links along the 00 direction (both in the direct and in the dual lattice) intersect the cut, and therefore are shared between different replicas. As a consequence, a plaquette in the dual lattice that lies along the 00 direction has some links in replica $1$ and some in replica $2$. Imposing eq. (35) as a solution of the constraint in this geometry, we obtain the following equations
@@
(and similar relations for other replicas), i.e., if the gauge variables lying on the entangling surface are the same in the different replicas. Also in this case we obtain the same picture as the $\mathbb{Z}_{2}$ gauge theory: gauge fields on the conical singularity belongs to all replicas, and they belong to $4n$ different plaquettes.

### v2

@@ Abstract
The study of entanglement in gauge theories is expected to provide insights into many fundamental phenomena, including confinement. However, calculations of quantities related to entanglement in gauge theories are limited by ambiguities that stem from the non-factorizability of the Hilbert space. In this work we study lattice gauge theories that admit a dual description in terms of spin models, for which the replica trick and Rényi entropies are well defined. In the first part of this work, we explicitly perform the duality transformation in a replica geometry, deriving the structure of a replica space for a gauge theory. Then, in the second part, we calculate, by means of Monte Carlo simulations, the entropic c-function of the $\mathbb{Z}_{2}$ gauge theory in three spacetime dimensions, exploiting its dual description in terms of the three-dimensional Ising model.
@@ 1 Introduction
In quantum field theories the entanglement entropy is affected by ultraviolet divergences: for highly excited states the divergence is proportional to the spatial volume of the system Bianchi:2021aui , whereas ground states exhibit the so-called “area law” Eisert:2008ur , namely the ultraviolet divergence is proportional to the area of the entangling surface $\partial A$ separating $A$ and $B$. Therefore, a proper regularization is required, in order to identify the subleading contributions, which encode important physical information about the “resolution scale” at which the degrees of freedom of the system are probed Casini:2010kt ; Liu:2012eea ; Casini:2012ei ; Myers:2012ed ; Klebanov:2012va ; Liu:2013una ; Jokela:2024cxb , the geometry of the system Klebanov:2012yf ; Casini:2014yca ; Mezei:2014zla ; Faulkner:2015csl ; Bianchi:2016xvf and the nature of the state itself Casini:2008cr . For a system defined in $D=d+1$ spacetime dimensions, it is then convenient to consider a geometry in which the entangling surface does not depend on the size of the subsystem $A$; a common choice is the slab geometry, in which $A$ is maximally extended in all space-like directions except for the one separating $A$ and $B$, where it has a finite size $l$. One can then define the ultraviolet-finite entropic (Rényi) c-function as Nishioka:2006gr ; Casini:2006es ; Casini:2004bw
@@
Calculations of such highly non-local quantities are typically challenging, both analytically and numerically. A tool which is commonly used is the replica trick Calabrese:2004eu ; Calabrese:2009qy , that allows one to express the trace of a power of the reduced density matrix appearing in eq. (1) in terms of ratios of partition functions,
@@
A non-trivial aspect in the study of entanglement entropy in gauge theories, however, is the ambiguity of the definition itself in the presence of a local symmetry. Gauß’s law and consequently the non-local constraint defining physical states makes the physical Hilbert space of a gauge theory non-factorizable. This problem has been extensively discussed in the literature, and different definitions of entanglement entropy have been proposed Buividovich:2008gq ; Donnelly:2011hn ; Agon:2013iva ; Casini:2013rba ; Radicevic:2014kqa ; Ohmori:2014eia ; Donnelly:2014gva ; Ghosh:2015iwa ; Aoki:2015bsa ; Chen:2015kfa ; Radicevic:2015sza ; Soni:2015yga ; VanAcoleyen:2015ccp ; Radicevic:2016tlt ; Soni:2016ogt ; Aoki:2016lma ; Lin:2018bud , among which two approaches are most commonly used: one can either extend the Hilbert space of states, including gauge non-invariant ones to ensure factorizability; or work at the level of the algebra of observables defined over a specific region, specifying which generators have to be included in region $A$ and which ones in its complement. It turns out that for some choices of the algebras the two approaches are equivalent Soni:2015yga , and lead to the following general result: the trace which defines the entanglement entropy splits into the sum of different contributions, labeled by the eigenvalues of the operators at the boundary between the system $A$ and its complement, which in general belong to the center of the algebra of observables of $A$. As a result, in Abelian gauge theories defined on the lattice the entanglement entropy admits the following decomposition
@@
In this work we address one facet of this problem, focusing on the construction of a replica space for gauge theories through a duality transformation. For gauge theories defined in three spacetime dimensions, the latter can be carried out through a Kramers–Wannier transform Kramers:1941kn ; Kramers:1941zz , which maps the gauge theory to a spin system. It is well known that such dualization is not straightforward when the gauge theory is defined on a space with a boundary and there exist operators (including, in particular, those defining Gauß’s law) defined on the boundary. However, it was shown in different works Casini:2014aia ; Radicevic:2016tlt ; Moitra:2018lxn ; Lin:2018bud that, even though the entanglement entropy is not preserved by the duality transformation, its universal part, encoded in the entropic c-function, is exactly mapped in the continuum limit. Note that the investigation of quantum lattice models by means of duality was recently addressed (in the Hamiltonian formalism) in ref. Fazza:2022fss , albeit with a focus on aspects different from the ones of the present work.
In the first part of this work, we explicitly perform the duality transformation starting from a spin model defined on a lattice in three spacetime dimensions with an arbitrary positive integer number of replicas, to derive the corresponding geometry for the dual gauge theory. As we discuss in what follows, the physical results we obtain from our construction are purely geometric, namely they do not depend on the degrees of freedom of the theory (nor on the details of how they are assigned to $A$ or $B$); rather, the duality transformation gives us a well-defined prescription on how to implement a conical singularity on the lattice. The logical scheme of how the direct and the dual formulation of a theory are connected with each other, in the Hamiltonian and in the Lagrangian formalisms, is summarized in figure 1 (which also includes the references to previous literature, in which the various aspects discussed herein were studied).
@@
The structure of the article is the following. In section 2 we review the generalities about the duality transformation, and discuss its formulation in the presence of replicas; we mostly focus on the case of the Ising model, considering both the case of two and three spacetime dimensions, but we also discuss the extension of our results to other gauge theories, including $\mathrm{U}(1)$ and $\mathbb{Z}_{N}$ gauge theories (and briefly comment on the generalization to higher dimensions and to non-Abelian gauge theories). Section 3 illustrates the connection between the Hamiltonian and the Lagrangian formulations in the presence of replicas, with a detailed discussion of the $(1+1)$-dimensional Ising model case study, while section 4 is devoted to a Monte Carlo study of the entropic c-function of the three-dimensional $\mathbb{Z}_{2}$ gauge theory. Finally, in section 5 we discuss our findings and comment on possible generalizations thereof. The appendix contains a discussion of the rôle of boundary conditions in the duality transformation (section A), a study of the dual replica space for other two-dimensional spin models (section B), a derivation of the spin theory dual to the three-dimensional $\mathrm{U}(1)$ gauge theory (section C), and some technical details about our numerical simulations (section D).
@@ 2.2 Duality in the presence of replicas
Figure 2: Two-dimensional replica space. The vertical direction is the Euclidean-time direction, while the horizontal one is the spatial one. The red wavy line represents the cut; green links connect spins in different replicas, therefore the periodicity of the lattice in the subsystem $A$ is $n$ times the periodicity in the subsystem $B$.
@@
The geometry we are interested in is the one of $n$ replicas connected through a cut, and let us first consider the two-dimensional case, as sketched in figure 2. In a replica geometry, the expression (8) still holds, since its derivation is only based on the nearest-neighbor nature of the interaction and is independent from the geometry of the lattice. The difference in the duality transformation is in the representation of the $k_{i,\mu}$ variables that satisfy the constraint; clearly, the introduction of a cut does not affect the duality transformation for links not crossed by the cut, since the transformation only involves nearest-neighbor spins in the dual lattice and the links in the direct lattice separating them. Moreover, the duality transformation remains unchanged also for links crossed by the cut that are far from the cut boundary. [^1] This is related to the fact that the exact location of the cut is irrelevant. The only difference arises at the edges of the cut; in particular, in two dimensions the dual lattice has sites exactly at (the lattice analogue of) the branch points of the replica space. Consider, for example, the cartoon in figure 3, where two of the $n$ replicas are depicted: the equation expressing the constraint associated with the site $i$ of the direct lattice, in the replica labeled $1$, is
@@
Figure 4: Duality transformation in a three-dimensional replica space. The solid line is the entangling surface along the links of the dual lattice, the red region is the cut, the red dot the intersection between the link between the sites $i$ and $i+\hat{0}$ and the cut.
It is straightforward to carry out the dualization in the presence of replicas for the three-dimensional Ising model, too. In this case (see the sketch in figure 4) the “cut” is actually a surface, and the conical singularity is a line where the cut ends. As for the two-dimensional model, the way the duality maps degrees of freedom in the direct and in the dual formulation is not affected by the presence of conical singularities, for spins far enough from them. Consider, for example, the site $i$ in the direct lattice shown in figure 4; in the first replica, the Kronecker-delta constraint is of the form
@@
The presence of gauge fields shared among different replicas leads to some constraints on the gauge transformations that leave the physical properties of the system invariant. In particular, there exist plaquette terms that can be schematically written as $U_{\mbox{\tiny{edge}}}U^{(r)}_{\mbox{\tiny{staple}}}$, where $U_{\mbox{\tiny{edge}}}$ is the gauge field on the conical singularity and $U^{(r)}_{\mbox{\tiny{staple}}}$ is a staple in replica $r$ forming a plaquette together with $U_{\mbox{\tiny{edge}}}$. The presence of such terms require that gauge transformations performed on the sites along the entangling surface to act on all replicas. More precisely, a gauge transform on a site on the entangling surface acts upon $(4n+2)$ gauge fields: the $4n$ links orthogonal to the conical singularity attached to that point, and the two links along the entangling surface shared by all replicas. This restriction on the possible gauge transformations in a replica space was also noted in the context of $\mathrm{SU}(N)$ gauge theories Soni:2016ogt ; Jokela:2022fvh .
@@
The difference between the Rényi entropies in the gauge and in the spin theories is the presence of an ultraviolet-divergent term proportional to the length of the entangling surface, therefore contributing to the area law. This term makes the Rényi entropy of the gauge theory larger than the one of the Ising model, as was already pointed out in the literature Tagliacozzo:2010vk . The origin of this difference stems from the constraint on the number of possible gauge transforms that can be defined in a system with replicas, and results into a multiplicative constant affecting the partition function for the dual gauge theory, which obviously does not carry any physical information. Indeed, one can also note that the term $|\partial A|-1$ depends on the gauge choice at the entangling surface, as was also discussed in refs. Casini:2013rba ; Aoki:2015bsa ; Aoki:2016lma . Like in the $D=2$ case, physical information can be obtained from a suitably defined entropic c-function, in which the area-law term cancels out; in $D=3$ if the length of the entangling surface does not depend on the size of the region $A$, as in the slab geometry, the entropic c-function can be simply defined according to eq. (2), and one gets
@@
It is also worth noting that, in the opposite limit ($\beta\rightarrow\infty$, or $\beta^{\star}\rightarrow 0$) the result one would naïvely obtain is not the same as in the Hamiltonian formalism:
@@
So far, we assumed the entangling surface to lie on the lattice dual to the Ising-model lattice, that is, to be placed along the links of the dual gauge theory. However, it is also interesting to analyze a geometry in which the entangling surface lies on the links of the spin model; in this case there are spin degrees of freedom on the conical singularity. From the discussion of the $D=2$ case, we know that such degrees of freedom are shared among all replicas, and have a larger number of nearest neighbors, which here is $(4n+2)$, of which $4n$ in the directions orthogonal to the entangling surface and $2$ along the conical singularity.
@@
where $U_{\Box\times n}$ denotes the closed loop of length $4n$ that winds around the conical singularity. Note that the presence of two such loops of length $4n$ is necessary to satisfy the Kronecker-delta constraint around a point of the direct lattice with an enhanced number of nearest neighbors: for the $4n$ links touching a site on the conical singularity one can make the usual duality transformation according to eq. (12), but then there are only two link variables to be adjusted, such that in the sum $\sum_{i}k_{i,\mu}$ all gauge variables appear twice. As a consequence, the two $k$ variables along the conical singularity must be built from a larger number of gauge variables. Note that this geometry was already discussed in ref. Chen:2015kfa , whose authors carried out a strong-coupling calculation of the Rényi entropy in this “central-plaquette” geometry, where the central plaquettes are the ones encircling the conical singularity, as sketched in figure 5.
@@
Figure 5: Central-plaquette geometry. The blue loop crosses the cut (pink surface) and therefore has length $4n$, where $n$ is the number of replicas.
@@ 2.3 $\mathrm{U}(1)$ and $\mathbb{Z}_{N}$ models
In the presence of replicas, the duality transformation can be generalized along the lines that we discussed for the three-dimensional Ising model; the geometry we consider is the same as in figure 4, but with $\phi$ variables instead of $\sigma$ ones. The constraint associated to the site $i$ of the left replica is again expressed by eq. (19). Note that, with the cut placed as in figure 4, all links along the 00 direction (both in the direct and in the dual lattice) intersect the cut, and therefore are shared between different replicas. As a consequence, a plaquette in the dual lattice that lies along the 00 direction has some links in replica $1$ and some in replica $2$. Imposing eq. (35) as a solution of the constraint in this geometry, we obtain the following equations
@@
(and similar relations for other replicas), i.e., if the gauge variables lying on the entangling surface are the same in the different replicas. Also in this case we obtain the same picture as in the $\mathbb{Z}_{2}$ gauge theory: gauge fields on the conical singularity belongs to all replicas, and they belong to $4n$ different plaquettes.